OpenStudy (anonymous):

Prove that ω^248+ω^154+5i^400=4 Note: (i=√-1 and ω is the complex cube root of unity)

hartnn (hartnn):

whats w^248? can u calculate that?? knowing w^3=1???

OpenStudy (anonymous):

i think it will break into parts (ω^248)

hartnn (hartnn):

....tell me whats w^6?

OpenStudy (anonymous):

sorry,i did'nt get you

hartnn (hartnn):

okk,as w^3 = 1 w^6=w^3*w^3=1*1=1 so any power of w which is multiple of 3 must equal 1 !!! so,w^246=1 so w^248=w^246*w^2=w^2 got this???

OpenStudy (anonymous):

there is an example on breaking in my book: ω^32 = (ω^3)^10.ω^2 can i do this by above example

hartnn (hartnn):

yes,thats what i did,i break w^248 as (w^3)^82.w^2 to get w^2 only u got that ???

OpenStudy (anonymous):

ok

hartnn (hartnn):

so,can u do same thing for w^154??

OpenStudy (anonymous):

yes

hartnn (hartnn):

then tell me what w^154 equals?

OpenStudy (anonymous):

ω^154=ω^153.ω (ω^3)^51.ω ω

OpenStudy (anonymous):

sorry i am slow in typing

hartnn (hartnn):

correct,so now your 1st 2 terms are w+w^2 to find that from 1+w+w^2=0 is easy....

OpenStudy (anonymous):

but how will 5i^400 be solved

OpenStudy (anonymous):

i^400 = (i^4)^100 = 1

OpenStudy (anonymous):

ω^248+ω^154+5i^400 = ω + ω^2 + 5(1) = ω + ω^2 + 1 + 4 = 1+ω + ω^2 + 4 = ...

OpenStudy (anonymous):

5i^400= 5(-1)^400 = 5 so, w^2+w+5 -1+5=4 am i right

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

5i^400 = 5(√-1)^400

OpenStudy (anonymous):

how will 5(√-1)^400 be solved

hartnn (hartnn):

i=swrt(-1) so i^2=1 so i^400=(i^2)^200=1^200=1

OpenStudy (anonymous):

i= \(\sqrt{-1}\) \(i^2 = (\sqrt{-1})^2 = -1\) \(i^3 = (\sqrt{-1})^3= -1i = -i\) \(i^4 = (\sqrt{-1})^4=(-1)^2 = 1\)

OpenStudy (anonymous):

i^2 is -1....

hartnn (hartnn):

sorry, i=sqrt(-1) so i^2=-1 so i^4=1 so i^400=(i^4)^100=1^100=1

OpenStudy (anonymous):

what?? w^2+w+1=0 we have to prove that it equals 4

hartnn (hartnn):

we already did,1st 2 terms = w+w^2 =-1 3rd term =5 i^400=5*(i^4)^100=5*(1^100)=5 5-1=4

OpenStudy (anonymous):

oh i forgot about 5 :D thanks

OpenStudy (anonymous):

5*(i^4)^100=5*(1^100) how did i^4 become 1

OpenStudy (anonymous):

I did explain above. i^4 = (i^2)^2 = (-1)^2 = 1 i^2 = \(\sqrt{-1} \times \sqrt{-1} = -1\)